A cubicle block of wood of edge 0.2m weighing 6.4kgf floats on water. Find the depth of immersion. Given, density of water 1000kg per cubic metre. Show working out please. 

Grade:12

2 Answers

Arun
4 years ago
s = 30 cm = 0.3 m
m = 21.6 kg
# Solution-
Let h be depth of immersion and Vi bs volume of immersed part.
Here, liquid displaced is equal to weight of wooden block.
mg = Vi.ρ0.g
m = s×s×h.ρ0
6.4 = 0.2^2 × 1000 × h
h = 0.16 m
h = 16 cm
 
Vikas TU
4 years ago
Dear student 
30 × 30 × 10^-4 × h × g = 21.6
Let h = the depth of immersion then
H= 24
Good Luck 
Cheers